19. -/1 POINTS TANAPCALC10 6.R.054. Find the area of the region under the curve y = 3x2 + 6x + 5 from x = -1 to x = 2.
Answer #1

At first we draw the region bounded by the curve and linens. Then evaluate the area of the region by using definite integral. Given cuore y= 3x²+60+ 5 > y = 32+62+ 3+2 > y = 3(x²+2x+1)+2 y-2= 3 (2+1) This is the equation of the parabola with verster (